Z^2dx+(z^2-2yz)dy+(2y^2-yz-xz) dz=0
1
Expert's answer
2020-12-28T13:30:21-0500

Comparing the given equation with the standard equation "Pdx+Qdy+Rdz=0," we get

"P=z^2,Q=z^2-2yz,R=2y^2-yz-xz"

The given equation is homogeneous of degree 2. Now first of all, we test the condition of integrability.

"P\left(\frac{\partial R}{\partial y}-\frac{\partial Q}{\partial z}\right)-Q\left(\frac{\partial R}{\partial x}-\frac{\partial P}{\partial z}\right)+R\left(\frac{\partial Q}{\partial x}-\frac{\partial P}{\partial y}\right)\\
=z^2(4y-z-2z+2y)-(z^2-2yz)(-z-2z)+(2y^2-yz-xz)(0-0)\\
=6yz^2-3z^3+3z^3-6yz^2=0"

Hence, the equation is integrable.


Substitute "x=uz,y=vz" into the equation.

This implies that, "dx=udz+zdu,dy=vdz+zdv"

"z^2(udz+zdu)+z^2(1-2v)(vdz+zdv)+z^2(2v^2-v-u)dz=0"

With this, the coefficient of "dz" will be 0. SO, we have;

"z^3du+z^3(1-2v)dv=0"

Divide through by "z^3"

"du+(1-2v)dv=0"

Integrate through

"u+v-v^2=C"

But, "u=\frac{x}{z},y=\frac{y}{z}"

"\frac{x}{z}+\frac{y}{x}-\frac{y^2}{z^2}=C\\
xz+yz-y^2=Cz^2"


Hence, the general solution of the equation is "xz+yz-y^2=Cz^2" .